OpenStudy (anonymous):

Will fan and give medals please help!!!! Evaluate the function , when g(t)=5/2t-3/2 t=3

OpenStudy (anonymous):

A. 6 B. 9 C. 12 D. 13.5 I really need someone to explain!!!

OpenStudy (abb0t):

This is verys traight forward, you see how it says \(\sf \color{red}{t=3}\), well all you need to do is plug in \(\sf \color{red}{3}\) whereever you see a \(t\) in the equation given.

OpenStudy (anonymous):

ok but how will I solve with the fractions?

OpenStudy (abb0t):

Just plug everything in.

OpenStudy (anonymous):

and what about g(t) what does that mean?

OpenStudy (abb0t):

g(t) means that \(g\) is a function composed of \(t\).

OpenStudy (anonymous):

So what do I do with that?

OpenStudy (abb0t):

You solve for it. Which is what you're doing. It is the same thing as saying g(3). Because t=3, remember?

OpenStudy (anonymous):

so then it ends up being 3g?

OpenStudy (abb0t):

For example, if you have g(t) = \(\sf \frac{11}{3t}\) and t=11, that means you have \(\sf \large \frac{11}{3\color{red}{(11)}}\) = \(\sf \frac{11}{33}= \frac{1}{3}\). Does that make sense?

OpenStudy (anonymous):

yeah kinda can you check my answer?

OpenStudy (abb0t):

3g is \(\color{red}{NOT}\) the answer.

OpenStudy (anonymous):

Right its 6 right?

OpenStudy (abb0t):

Let's check

OpenStudy (abb0t):

\(\sf \frac{5}{2}(3) - \frac{3}{2} = \frac{15}{2}-\frac{3}{2}=\frac{12}{2} = \color{red}{6}\)

OpenStudy (anonymous):

So I was right?

OpenStudy (abb0t):

YES!

OpenStudy (anonymous):

Thanks a million times could you help me on another one if I tag you in the question?